OpenStudy (anonymous):

f(x)=-x2-3x+5, find -5. (x2=xsquared)

jimthompson5910 (jim_thompson5910):

I'm assuming you mean f(-5) f(x)=-x^2-3x+5 f(-5)=-(-5)^2-3(-5)+5 f(-5)=-25-3(-5)+5 f(-5)=-25+15+5 f(-5) = -5

OpenStudy (anonymous):

yes that is what i got thank you! 

OpenStudy (anonymous):

f(-5) = -5